Simplify (9.0×10^9)[20×10^-8/d+0.1)-5.0×10^-8/d^2)=0


1
Expert's answer
2021-04-15T07:27:18-0400

"9x10^9(\frac{20x10^{-8}}{d}-\frac{5x10^{-8}}{d^2)}=0"

"9x10^9(\frac{20x10^{-8}d-5x10^{-8}}{d^2})=0"

"9x10^9(\frac{5x10^{-8}(4d-1)}{d^2})=0"

"(\frac{450(4d-1)}{d^2})=0"

"4d-1=0"

d=0.25